OpenStudy (anonymous):

can somebody please help me? 5/8=x+6 solve for x

OpenStudy (anonymous):

that's not the answer

OpenStudy (akashdeepdeb):

So just subtract 6 from both sides :) You get! = 5/8-6 = x Or x = -43/8!! :D Get it? :D

OpenStudy (anonymous):

thank you!!!

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

can you help with another?

OpenStudy (akashdeepdeb):

btw Answer is NOT important! What's important is iif you get the concept/ method! Okay? :)

OpenStudy (anonymous):

U+3/8=-3/4 Solve for U

OpenStudy (anonymous):

if you could explain to me thank you. ive been stuck all day doing this

OpenStudy (akashdeepdeb):

Just do the same thing U = -3/4-3/8 Answer= -9/8 Getting why and how? :D

OpenStudy (anonymous):

I think so.

OpenStudy (akashdeepdeb):

See for aolving simultaneous equations! All you need to do is! Bring the VARIABLY to one side and bring the constant numbers to the other and THEN calculate!! :D

OpenStudy (anonymous):

-2/5+V=1/4 so I add -2/5 to each side?

OpenStudy (anonymous):

but don't I have to find a common denominator?

OpenStudy (akashdeepdeb):

NO! See you want to get RID of the constant on the LHS so Add 2/5 :D Okay? :D

OpenStudy (akashdeepdeb):

After that find the LCM! Like 1/4+2/5 = 1*5/4*5 + 2*4/5*4 [NOW YOU HAVE COMMON DENOMINATOR 20] (5+8)/20 = 13/20!! Getting me? :D

OpenStudy (anonymous):

okay. one last one let me see if im doing it right. its -6=x-3/5 add -3/5 to each side?

OpenStudy (akashdeepdeb):

Kinda...... ADD +3/5 to EACH side to get rid of that -3/5 Because? -3/5 and +3/5 cancel each other!! :D

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

so then... now im confused again

OpenStudy (anonymous):

-6+-3/5?